# Shallow learning enables real-time inference of molecular composition from spectroscopy of brain tissue

### v3

@@ 3 Results
3.1 Data.
For our experiments, we used two types of modified Beer-Lambert law formulation, with and without scattering, on two types of spectral datasets: broadband NIRS data for which the spectra were measured in light transmission mode [15], and hyperspectral data which were obtained in reflection mode [16].
Broadband NIRS.
The first dataset is composed of broadband NIRS spectra from a study analyzing 27 piglets’ brains in which a hypoxia state was induced [15]. The piglets were monitored for several hours, during which the carotid arteries were surgically isolated. This allowed controlling the blood flow to the brain. The details of the intervention protocol are described in [15]. The measurements contain around eight thousand spectra per piglet, of which we use the first thousand measurements (corresponding to the first $\sim$2.5 hours of measurement) where HI and recovery after HI are visible. As Eq. 2 requires defining a reference spectrum, analogously to [15], we used a spectrum at the very beginning of optical monitoring (i.e., before HI) for reference. We normalized the spectra with respect to dark noise and white reference. The normalized NIRS spectra before and after the intervention inducing hypoxia are shown in Fig. 3 (left), and predictions of the concentrations change over the course of the optical monitoring are shown in Fig. 3 (right). Out of the 27 piglets in the dataset, 25 had data available during HI, such that 19 were used for training, two for validation, and four for testing. For this dataset, we predict three types of molecules: oxyhemoglobin, deoxyhemoglobin, and differential cytochrome-c-oxidase (CCO) [^3] As the total CCO concentration may be assumed to not change within a few hours, the oxidized-reduced difference spectrum $\mu_{a}^{\text{diffCCO}}$ may be used to infer changes of both oxidized and reduced CCO [29]. , i.e. $\{\delta c_{i}\}=\{\delta c_{HbO_{2}},\delta c_{HHb},\delta c_{\text{diffCCO}}\}$, where $\delta c_{\text{diffCCO}}=\delta c_{oxCCO}-\delta c_{redCCO}$. We neglected the potential contribution to the spectra from water and fat due to their minimal change in concentrations during the monitoring and rather flat absorption spectra in the considered range of wavelengths between 740 and 900 nm [15, 30].
Figure 3: Optical spectra from the broadband NIRS study in [15] before and after inducing hypoxia in the piglet’s brain (left). On the right, predictions of the molecular concentration change $\{\delta c_{HbO_{2}},\delta c_{HHb},\delta c_{\text{diffCCO}}\}$ over the course of the optical monitoring (for the Beer-Lamber model without scattering).
Hyperspectral data.
The second dataset we used consists of hyperspectral data from the Helicoid project [16]. The Helicoid dataset comprises brain HSI images obtained in surgical conditions from 22 patients diagnosed with glioma. The HSI images provide a high spectral resolution of 826 bands spread between 400 and 1000 nm and a 2D spatial resolution of a few hundred pixels in each dimension. The images were also expert-annotated into three tissue classes: normal and tumor tissues, as well as blood vessels. Typical hyperspectral image and corresponding spectra are shown in Fig. 4. Different from the NIRS dataset, as a reference spectrum, we used a spectrum of the pixel belonging to the blood vessel class. We then subtracted the reference spectrum from all other spectra in the image. Besides predicting oxyhemoglobin $\delta c_{HbO_{2}}$ and deoxyhemoglobin $\delta c_{HHb}$, we again infer the differential cytochrome-c-oxidase concentration due to its role in capturing oxidative metabolic activity. We separately predicted oxidized cytochrome-c-oxidase and reduced cytochrome-c-oxidase, as the total CCO concentration may not be assumed to remain constant in space. We also predict water and fat since, for these molecules, one cannot assume minimal concentration change across different tissue types as in the case of the NIRS spectra. Fig. 5 showcases examples of molecular inference for the HSI images from the Helicoid dataset. Out of the nine patients with glioblastoma in the dataset, six with distinct class labeling were chosen, and three patients were used for training, one for validation, and two for testing. Note that patients might have multiple images taken, such that the training set consists of five, the validation set of one, and the test set of three images.
Figure 4: Optical spectra from the HSI study of patients diagnosed with glioma [16] for different tissue types: tumor, normal tissue, and blood vessels.
Figure 5: An example of HSI image from the Helicoid dataset [16] (left). Each pixel in the shown 2D image possesses a spectral signature with 826 bands. From this signature, we predict the molecular concentration change for hemodynamic $\delta c_{HbT}=\delta c_{HbO_{2}}+\delta c_{HHb}$ (middle), and metabolic $\delta c_{\text{diffCCO}}=\delta c_{oxCCO}-\delta c_{redCCO}$ characterization (right).
3.2 Experiments.
Scattering vs Non-scattering.
First, we test different Beer-Lambert formulations - with and without scattering. This test is motivated by our desire to assess whether networks can substitute not only linear solvers (without scattering) but also non-linear ones (with scattering) since the latter require a large computation time for optimization.
@@
Figure 6: Comparison between predictions using linear (no scattering) and non-linear (with scattering) models for NIRS (left) and HSI (right) spectra. The inclusion of scattering into the formulation of the Beer-Lambert law notably improves the spectral fit to real data.
@@
Fig. 6 shows spectral fits using both formulations. The model with scattering provides a clearly better fit. It allows us to better describe the peak around 760 nm for the NIRS data, while for the HSI data, the inclusion of scattering is often merely necessary for an accurate fit of the spectra. The necessity of the scattering consideration in the Beer-Lambert model for the HSI data can be explained by the more pronounced contribution of the scattering process  [^4] For the HSI data, we infer the difference in molecular composition between different spatial locations on an image, i.e., between different tissue types. The scattering property across brain tissues can significantly vary, and thus, the scattering shapes markedly the differential spectra. In contrast, for the NIRS data, we perform the differential spectroscopy analysis not in space but in time (comparing two spectra for the same location taken at different time points), meaning that the molecular inference is performed for the same tissue type..
Note, as a by-product of our analysis, we observe that the molecular inference driven by the physical model leads to better contrasting the blood vessel tree compared to the RGB image, Fig. 5 (middle), and capturing metabolic activity (right).
Synthetic vs Real.
i
ii
@@
Figure 7: Comparison between inference of the molecular composition using the standard optimization methods and proposed network-based inference for training strategy (a) and (b) on both datasets. The top row compares strategy (a), left, with strategy (b), right, for the NIRS dataset. The bottom row compares the inference of the hemodynamic signal of the optimization-based result, left, with strategy (a), middle, and strategy (b), right. Despite the fact that the training distribution is different from the test distribution (especially for strategy (a) with synthetic data used for training), the network predictions are closely aligned with the predictions by the standard solver.
Fig. 7 demonstrates the results of the experiment in which we test the network trained on synthetic data collected according to strategy (a) and on real data according to strategy (b). The latter provides qualitatively closer fits for the shown piglet, which is also quantitatively supported as presented in table 1. Nevertheless, the predictions from both strategies are close to the ground truth predictions obtained by the least-squares optimizer, especially for NIRS data. Intuitively, this behavior is not within expectations as the spectra of synthetic data on which the network was trained and real data on which it was evaluated notably vary (e.g., due to noise). We attribute such behavior to the fact that the main spectral feature that networks learn to focus on is the global functional shape of the spectra. The local behavior of the spectral function is less informative, and thus, the local perturbations, i.e., random noise, do not affect the network performance significantly. This reasoning is further reinforced when comparing the network performance between NIRS and HSI data. For the HSI data, the noise has a pronounced wavelength dependency, in contrast to NIRS data (see Fig. 4), which results in the network performance degrading more notably compared to NIRS data.
@@
Table 1: Quantitative performance comparison of the different Beer-Lambert models and network training strategies on the test set of the two spectral datasets. To compare the two (non-scattering and scattering) models, we compute the mean absolute error of the spectral fit (denominated as ’Spectral MAE’) between the ground truth observed and predicted signals. The comparison of the two network training strategies is performed by assessing the mean absolute error of each strategy between the network and optimization-inferred concentrations (denominated as ’Concentration MAE’) of all considered chromophores. In the case of the HELICoiD dataset, only pixels labeled as normal, tumor, or blood were considered for these computations. The best-performing model and strategy for each individual is highlighted in bold.
Another explanation for this behavior can be that here, we try to learn with a neural network a solution to a linear system (or quasi-linear system, in the case of the model with scattering, where scattering contribution to the reflection spectra is minor compared to the absorption, as for NIRS data). We hypothesize that upon network training on the dataset representing solutions to linear systems, the network weights are learned to minimize the contribution from non-linear network units (e.g., activation functions), as such units are unnecessary to establish a linear mapping. Generally speaking, the non-linear units are a virtue and a curse of the networks. The virtue, because they allow to learn an arbitrary mapping, and the curse, because it is the very sequential application of non-linear units that blows up the predictions for input data unseen during training. But if the contribution from non-linear units of the network is minimized, we would not expect such blow-ups of the network’s predictions.
Within expectation were, however, the results obtained by the strategy (b). In contrast to training on the synthetic data, strategy (b) allows us to infer molecular composition almost identical to the convex optimization solver for both NIRS and HSI spectra.
Computational time.
@@
Importantly, the proposed network-based optimization comes with a significant speed-up in computational time. In Fig. 8, we show a comparative analysis for performing chromophore composition inference using standard least-squares solvers (based on gradient update or pseudoinverse) and our proposed approach. As solving the linear system using the pseudoinverse requires the least amount of matrix multiplication, this method provides the fastest compute.
@@
However, with the growth of the number of spectra for which we solve the optimization task, the matrix size for the inversion increases, and thus the computational time increases. Starting from ca. $10^{4}$ amount of spectra, the proposed network having a fixed amount of computational units becomes superior in terms of optimization time. Such runtime will remain approximately constant with a further increasing number of spectra, assuming sufficient GPU memory is available. More importantly, for non-linear systems, which are here represented as a Beer-Lambert model with the inclusion of scattering, one cannot utilize the pseudoinverse and has to resort to non-linear solvers like the ones based on gradient update. Such solvers are two to three orders of magnitude slower than the neural network approach, which has fixed compute time for linear and non-linear systems. As Fig. 8 shows, it takes ca. 0.7 ms for the network to infer biochemical composition for $10^{5}$ spectra on NVIDIA GeForce MX450 with 2048 MiB.
4 Conclusion
In the paper, we present proof of a data-driven concept for inferring molecular composition change from optical reflection spectra of brain tissue. We test the approach on various datasets (broadband NIRS and HSI data) and physical models (with and without scattering, i.e., linear and non-linear). We achieve a significant speed-up compared to traditional least-squares optimization, reaching subsecond time for inferring molecular composition. The subsecond timing stays the same for linear models (without scattering) and non-linear models (with scattering). Such timing would allow for real-time tissue characterization for both NIRS and HSI imaging.
@@ 4 Conclusion
In support of open science, all our code is available under https://github.com/HyperProbe/SpectraFit. The HSI data is publicly available by the HELICoiD project. The bNIRS data may be obtained upon request.

### v6

@@ 3 Results
3.1 Scattering vs Non-scattering.
First, before discussing the learnable methods for molecular inference, we test different Beer-Lambert law formulations - with and without scattering - to elucidate limits of applicability of both models. For the case of piglets undergoing HI, it is widely assumed that the 780nm to 900nm range is predominantly dominated by absorption, with scattering being only a minor contributor to the overall measured spectrum [38]. As measurements in the piglet dataset below the 780nm threshold were available, we opted to extend the model fitting range from 740nm to 900nm. This test is motivated by our desire to assess whether a linear model (without scattering) would still be sufficient to describe the broader spectroscopy measurement of brain tissue.
@@
Figure 6: Comparison between predictions using linear (no scattering) and non-linear (with scattering) models for bNIRS (left) and HSI (right) spectra. The ground truth (GT) attenuation is computed from the real spectra difference. The inclusion of scattering into the formulation of the Beer-Lambert law notably improves the spectral fit to real data.
Fig. 6 and 7 show spectral fits and inferred molecular concentrations using both formulations. The model with scattering provides a clearly better fit. It allows us to better describe the peak around 760 nm for the bNIRS data, while for the HSI data, the inclusion of scattering is often merely necessary for an accurate fit of the spectra in this wavelength range. This finding is consistent across the dataset, as shown in table 1.
Figure 7: On the left, comparison between chromophore predictions using the linear model without scattering (”Standard model”) and non-linear model with scattering (”Scattering model”) for the bNIRS dataset. On the right, the predicted coefficients $s$ and $b$ for the scattering term in Eq. 3. Note that the purple curve represents an overlap across the two red and blue curves.
Figure 8: Histograms showing mean relative improvement of the spectral fit MAE for presented scattering model, compared to the linear model, across all piglets in the dataset for all wavelengths (left), in the range 740nm-780nm (middle), and in the range 780nm-900nm (right) in the broadband NIRS dataset. The x-axis represents the relative improvement $r$ between the two models, and the y-axis shows the number of piglets that achieved the corresponding mean relative improvement $r$. The dashed line signifies improvements below zero, i.e. cases where the spectral fit worsened.
To show that the model with scattering can significantly improve model fits, especially for the higher frequency portion of the spectrum, we evaluate the relative improvement $r$ in terms of the spectral fit of the scattering model compared to the non-scattering model. We use mean absolute error (MAE) as a measure of the fit and compute it for all piglets across different spectrum bands.
The spectral fit MAE is calculated by
$\text{MAE}=\frac{\sum_{i=1}^{n}|\Delta A_{model}(\lambda_{i})-\Delta A_{data}(\lambda_{i})|}{n}$ (5)
where $\Delta A_{model}(\lambda_{i})$ represents the model-inferred attenuation at wavelength $\lambda_{i}$, and $\Delta A_{data}$ the real measured attenuation, respectively. The relative mean improvement is computed by
$r=\frac{\text{MAE}_{scatter}-\text{MAE}_{linear}}{\text{MAE}_{linear}}$
comparing the improvement in mean error between the scattering and linear models, computed across all timepoints.
The results of such computation for all piglets are shown in figure 8, where we observed a mean relative improvement of 15.7% over the full fitting range. Such improvement is especially noticeable in the 740nm to 780nm range, where the mean improvement of the distribution almost doubled at 30.8%. The spectral fit does not improve significantly in the 780nm to 900nm range, with the mean relative improvement of the distribution being at merely 5.6%. For one of the 25 piglets, we observed that the spectral fit slightly worsened in the 780nm to 900nm range through the nonlinear model. However, the fitting MAE is only worse by 0.6%, and the spectral fit was still better for the overall range and in the 740nm-780nm range. We therefore can still confidently conclude that the presented model is able to fit the piglets’ measured spectra more closely, especially for presumed scattering-dominated bands.
@@
The necessity of the scattering consideration in the Beer-Lambert model for the HSI data can be explained by the more pronounced contribution of the scattering process. For the HSI data, we infer the difference in molecular composition between different spatial locations on an image, i.e., between different tissue types. The scattering property across brain tissues can significantly vary, and thus, the scattering shapes markedly the differential spectra. In contrast, for the bNIRS data, we perform the differential spectroscopy analysis not in space but in time (comparing two spectra for the same location taken at different time points), meaning that the molecular inference is performed for the same tissue type.
In conclusion, we find that the non-linear model is especially helpful in describing scattering-dominated bands. However, the linear model may still be used when absorption is the prevalent physical effect.
Figure 9: Comparison between inference of the molecular composition using the standard optimization methods and proposed network-based inference for training strategy (a) and (b) on the bNIRS dataset. The top row compares both strategies when using the linear model, where highly accurate neural network predictions are visible in both cases. The bottom row compares the strategies when using the non-linear model, with strategy (b) delivering noticeably more accurate predictions.
@@
Figure 10: Comparison between inference of the molecular composition using the standard optimization methods (Opti) and proposed network-based inference (NN) for training strategy (a) and (b) on the Helicoid dataset. The top row compares strategy (a), left, with strategy (b), right, for the bNIRS dataset. The figure compares the inference of the hemodynamic signal of the optimization-based result, left, with strategy (a), middle, and strategy (b), right. Strategy (b) significantly improves results when using the non-linear model.
3.2 Evaluating different training strategies.
@@
Table 1: Quantitative performance comparison of the different Beer-Lambert models and network training strategies on the test set of the two spectral datasets. To compare the two (non-scattering and scattering) models, we compute the mean absolute error of the spectral fit (denoted as ’Spectral MAE’) between the ground truth observed and predicted signals. The two network training strategies are compared by assessing the mean absolute error of each strategy between the network and optimization-inferred concentrations (denoted as ’Concentration MAE’) of all considered chromophores. In the case of the HELICoiD dataset, only pixels labeled as normal, tumor, or blood were considered for these computations. The best-performing model and strategy for each individual is highlighted in bold.
Next, we evaluate the proposed machine learning approach in its ability to substitute both the linear absorption and the non-linear scattering model.
Fig. 9 demonstrates the results of the experiment in which we test the network trained on synthetic data collected according to strategy (a) and on real data according to strategy (b), for both linear and non-linear models. For the linear case, both strategies are able to correctly infer the concentrations. The solution to the linear model can merely be found by a matrix multiplication, i.e. the pseudoinverse, which is why both strategies are able to very accurately predict the optimization-inferred concentrations.
For the non-linear case, strategy (b) provides qualitatively closer fits. We also tested this model for the Helicoid dataset, where we found highly matching results by the use of strategy (b), as seen in Fig. 10.
These findings are also quantitatively supported by the results in Table 1.
3.3 Computational time.
@@
Importantly, the proposed network-based optimization comes with a significant speed-up in computational time. In Fig. 11, we show a comparative analysis for performing chromophore composition inference using standard least-squares solvers (based on gradient update or pseudoinverse) and our proposed approach. The used spectra for this comparative analysis are taken from the broadband NIRS dataset assuming the scattering model, i.e. they are in the 740nm-900nm range, with a total of 244 measured wavelengths per spectrum, and the underlying chromophores are oxyhemoglobin, deoxyhemoglobin, and differential cytochrome-c-oxidase.
As solving the linear system using the pseudoinverse requires the least amount of matrix multiplication, this method provides the fastest compute. However, with the growth of the number of spectra for which we solve the optimization task, the matrix size for the inversion increases, and thus the computational time increases. Starting from ca. $10^{4}$ amount of spectra, the proposed network having a fixed amount of computational units becomes superior in terms of optimization time. Such runtime will remain approximately constant with a further increasing number of spectra, assuming sufficient GPU memory is available. More importantly, for non-linear systems, which are here represented as a Beer-Lambert model with the inclusion of scattering, one cannot utilize the pseudoinverse and has to resort to non-linear solvers like the ones based on gradient update. Such solvers are two to three orders of magnitude slower than the neural network approach, which has fixed compute time for linear and non-linear systems. As Fig. 11 shows, it takes ca. 0.4 ms for the network to infer biochemical composition for $10^{5}$ spectra on NVIDIA GeForce MX450 with 2048 MiB. Overall, on our hardware, it takes between 2.5 and 3.1 seconds to run the neural network for one an image from the HELICoiD dataset (the largest among all tested data), from opening the normalized HSI image and loading the neural network into the GPU, to displaying the inferred concentrations.
@@
4 Discussion and Conclusion
First, in this paper we wanted to address the limitation of the existing machine learning approaches to infer molecular composition using a physical model. Predominantly, the training within these approaches is performed on synthetic data produced by the physical model, analogously to the strategy (a) above. However, real spectroscopy measurements include a few factors not considered by the modeling, such as instrumentation and physiological noise or other non-linear optical phenomena. To close the gap between synthetic and real data, several works propose incorporating various kinds of noise into the synthetic data [26]. But the realism of the used noise formulations and their sufficiency to close the gap can still be questioned. Instead, the learning scheme proposed in this paper results in training and testing carried out on the same type of data obtained by processing real spectra with traditional least-squares optimisation. Thus, no noise model is required to achieve accurate predictions.
However, we must admit that the predictions by the existing training approach are, nevertheless, close to the ground truth. Intuitively, this behavior is not within expectations as, again, the spectra of synthetic data on which the network was trained and real data on which it was evaluated notably vary. We attribute such behavior to the fact that the main spectral feature that networks learn to focus on is the global functional shape of the spectra. The local behavior of the spectral function is less informative, and thus, the local perturbations, i.e., random noise, do not affect the network performance significantly. This reasoning is further reinforced when comparing the network performance between bNIRS and HSI data. For the HSI data, the noise has a pronounced wavelength dependency, in contrast to bNIRS data (see Fig. 4), which results in the network performance (trained according to the existing approach) degrading more notably compared to bNIRS data.
Another explanation for this behavior can be that here, we try to learn with a neural network a solution to a linear system (or quasi-linear system, in the case of the model with scattering, where scattering contribution to the reflection spectra is minor compared to the absorption, as for bNIRS data). We hypothesize that upon network training on the dataset representing solutions to linear systems, the network weights are learned to minimize the contribution from non-linear network units (e.g., activation functions), as such units are unnecessary to establish a linear mapping. Generally speaking, the non-linear units are both a strength and a limitation of the networks. The strength is that they enable to learn an arbitrary mapping, while the limitation is the very sequential application of non-linear units that can cause the predictions to become highly unstable for input data unseen during training. But if the contribution from non-linear units of the network is minimized, we would not expect such instability of the network’s predictions.
Second, we want to understand the limits of applicability of the linear model (with no scattering) for describing brain tissue spectroscopy. Note that under the assumption of negligible contribution from scattering in the modified Beer-Lambert law, the Eq. 2 results in a linear system. Such a system can be efficiently solved using the pseudoinverse, achieving close to real-time computation. However, in general, when including non-linear terms in the Beer-Lambert law formulation (scattering, non-linear absorption effects, etc.), the system of equations can not be solved anymore using pseudoinverse. Our analysis reveals that including scattering can often be necessary to describe the spectroscopy measurements, especially for HSI data. But also for the bNIRS data, we have qualitatively observed that the nonlinear model seems to fit the wavelength range between 740nm and 780nm particularly well, Fig. 8. This finding was within expectations since the range between 780nm and 900nm is not regarded to be dominated by changes in scattering and has therefore been widely used in broadband NIRS technologies with concentrations inferred through the Beer-Lambert law solely considering absorption [38].
Importantly, as a by-product of our analysis, when we test the spectral unmixing on the brain tissue HSI dataset, we observe that the molecular inference driven by the physical model leads to better contrasting the blood vessel tree compared to the RGB image, Fig. 5 (middle), and capturing metabolic activity (right). Image semantics related to both hemodynamic and metabolic tissue properties could be highly valuable in assisting neurosurgeons during the process of tumor removal. Knowledge of the topology of the blood vessel tree would minimize the amount of undesirable resection of the vessels leading to bleeding. In turn, understanding the metabolic activity across the 2D field of view would allow for better separation of pathological tissue from the healthy parenchyma, reducing the time spent on intra- and post-operative biopsy.
To conclude, in the paper, we present a data-driven concept for inferring molecular composition change from diffuse spectroscopy of brain tissue. We test the approach on various datasets (bNIRS and HSI) and physical models of different complexity (with and without scattering, i.e., linear and non-linear systems). Importantly, we evaluate different training strategies for neural-networks-based molecular prediction. The proposed strategy provides predictions that are nearly identical to the traditional least-squares-fit method, making the learnable solver an accurate alternative. Finally, the method achieves subsecond time for simultaneous inference of molecular composition across a large number of spectra, allowing for real-time tissue characterization using bNIRS and HSI imaging modalities.
@@ 4 Discussion and Conclusion
In support of open science, all our code is available on github under https://github.com/HyperProbe/SpectraFit. The HSI data is publicly available by the HELICoiD project. The bNIRS data may be obtained upon request. Both datasets received ethical approval as described in the corresponding papers [38, 39].